OpenStudy (anonymous):

Algebra Question! What is the value of "b" in the triangle shown below? The triangle is a right triangle with an area of "24 inches ^2" the the height is "3b" and the length is "b". Answer Choices are: A.) –4 in. B.) 4 in. C.) ±4 in. D.) There is no solution. They never covered this in the book...

OpenStudy (anonymous):

B

OpenStudy (anonymous):

u know the area of tringle formula

OpenStudy (anonymous):

What type of correlation would you expect from a scatterplot of time you have traveled versus miles remaining until you reach your destination? strong negative strong positive weak positive weak negative

OpenStudy (anonymous):

1/2 x base x height = area 1/2 x b x 3b = 24 3b^2 = 48 b = +4 ,-4 but since its a length therefore +4 is the answer

OpenStudy (anonymous):

thats right